You have a credit card with an APR of 36%. The card requires a minimum monthly payment of 5% of the balance. You have a balance of $7000. You stop charging and make only the minimum monthly payment. What is the balance on the card after eight years?
Answer #1

The APR is 36% therefore monthly interest is = 36/12 = 3% .

A minimum payment of 5% is required every month.

The monthly change in balance= Minimum Payment % - Monthly Interest = 5% - 3% = 2%.

Decrease in balance every month will be 2%

Balance for month + 1 = Balance for month x 0.98

Total months in 8 years would be = 12 x 8 = 96 months

Balance after 8 years = $7000 x (0.98) ^ 96 = $7000 x 0.14378 = $1006.47 (rounded off)